OpenStudy (anonymous):

why the internal energy is 0 in case of isolated system? I mean, how the dq and dw are both 0

OpenStudy (anonymous):

internal energy is not zero, internal energy is a CONSTANT, hence the CHANGE in the internal energy is zero!!

OpenStudy (anonymous):

oh sorry ...i meant the same... how dW is 0

OpenStudy (anonymous):

well.. actually.. in an isolated system, i don't think the internal energy is a constant.. infact only dq is zero.. because no energy is allowed to flow in or out ( hence the term isolated), so there will be change in internal energy and it ll be used to do some work, so by using the first law of thermodynamics we have dq= dU + dW hence we put dq= 0 so dU= -dW.. and thus.. the work is done BY the system ( negative sign implies that).. an example of this would be an adiabatic process!

OpenStudy (anonymous):

u r right but u explained a closed system here...what abt isolated system??? in which dq is also 0 and dw=0 ... is it bcz of this..? dU=0 (bcz isloated system no heat supplied so, dq=0 and since, dU= dQ+dW so 0=0+dw => dw=0

OpenStudy (anonymous):

but my problem is how internal energy is 0 here

OpenStudy (anonymous):

internal energy is never zero.. i told you.. its a CHANGE that is zero.. !!

OpenStudy (anonymous):

If you consider an isolated system.. no energy is given to it, and no energy is given out from it as well.. hence how can it do work?

OpenStudy (anonymous):

may be its process is adiabatic,so W no 0